OpenStudy (anonymous):

Dr. Warren purchased some medical supplies for $670. Sales tax rate was 6.5%. How much did Dr. Warren spend on these supplies?

OpenStudy (anonymous):

670 • 6.5 = x

OpenStudy (anonymous):

x + y = answer :D

OpenStudy (anonymous):

No wait. Scratch that.

OpenStudy (anonymous):

670 ÷ 6.5 = 103.1 670 + 103.1 = x

OpenStudy (anonymous):

i got the first part of it whichis 670 * 6.5 which gives u 43.55. the theres another step

OpenStudy (anonymous):

you add 43.55 to the original 670. That makes the total purchase

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

Great job!

OpenStudy (anonymous):

thanks.